OpenStudy (anonymous):

(2xy+3y^2)/(2xy+x^2)=dy/dx y(1)=1 private solution ?

myininaya (myininaya):

without trying to myself we could try y=vx which means y'=v'x+v this could help or this may not help --- if that doesn't work, maybe we can see if it is exact

myininaya (myininaya):

actually first way seems to work

OpenStudy (calculusfunctions):

@rubcksy do you understand? @myininaya is correct because what you have is a homogenous equation of degree zero. Thus you can start with y = vx.

OpenStudy (anonymous):

y=vx dy=xdv+vdx i applied this transform but i cannot get the answers.maybe i have some mistakes at the other steps

OpenStudy (calculusfunctions):

First if\[y =vx\]then differentiating with respect to x yields\[\frac{ dy }{ dx}= x \frac{ dv }{ dx }+v\]